Evaluate the definite intergral "\int"01 18e3x+1dx (round to an interger).

1.         322

2.         311

3.         932

4.             965


1
Expert's answer
2021-07-25T11:06:54-0400

Let,

"3x+1=t"

differentiating above equation with respect to x

"3dx=dt"

"dx=\dfrac{dt}{3}"


When "x=0;\space t=1"

"x=1;\space t=4"


"\u222b_0^1 18e^{3x+1}dx"

Substituting "(3x+1=t)" in above problem and changing limits of integration

"= ∫_1^4 \dfrac{18e^{t}dt}{3}"

"=6\u222b_1^4 e^{t}dt"

"=6 |e^{t}|_1^4"

"=6(e^4-e)"

"=311.27\approx311"